How many L of a 45% sugar solution must be mixed with a 4 L of pure water to make a 27% solution?
Answer by mananth(16949)

------- percent ---------------- quantity
water 0 ---------------- 10 Liters
sugar solution 45 ---------------- x Liters
Total 27 ---------------- 10 + x Liters

0*10+45 *x=27 (10 + x )

45 x = 270 + 27 x

45 x - 27 x = 270 - 0

18 x = 270

/ 18

x = 15 Liters 45 % sugar solution